# Direct and inverse abstract Cauchy problems with fractional powers of almost sectorial operators

### v1

@@ Abstract.
We show the explicit solution operator of an abstract Cauchy problem involving a time-variable coefficient and a fractional power of an almost sectorial operator. Also, by using the representation of the solution operator, we solve the inverse abstract Cauchy problem for recovering the time-variable coefficient. As a complement, we also study similar problems by using almost sectorial operators that depend on a time-variable.
@@ 1. Introduction
In functional analysis, specifically in operator theory, the Hille–Yosida theorem characterizes the generators of strongly continuous semigroups of linear operators on Banach spaces. This characterization requests operators to be closed and densely defined. Moreover, the following estimate holds:
$\|(\lambda I-A)^{n}\|\leqslant\frac{C}{(\lambda-\omega)^{n}},\quad\lambda>\omega,$
@@
for all positive integers $n$ and any $\lambda$ in the resolvent set $\rho(A).$ Note that for contraction semigroups, the Hille–Yosida theorem can be easily characterized by those operators $A$ such that $\|(\lambda I-A)^{-1}\|\leqslant\frac{1}{\lambda},$ for any $\lambda\in\rho(A)$ with $\lambda>0.$ Nevertheless, there are different classes of important and interesting operators who do not satisfy the above estimates. Indeed, we recall the so-called almost sectorial operators, see Definition 2.1. More details about this type of operators can be found e.g. in [7, 14, 18]. It is known that operators satisfying condition (b) do not generate a $C_{0}$-semigroup since it does not satisfy the Hille-Yosida’s estimates. But, in this case, the operator $A$ will generate a type of semigroup called analytic semigroup of growth oder $\gamma$ (see Definition 2.2). This concept was first introduced by Da Prato in [5] for positive integer orders. The generalization for any positive order was given in different works e.g. [10, 11, 18, 20] and the references therein. These semigroups do not imply strong continuity at $t=0,$ and this is one of the main difference compare with the $C_{0}$-semigroups. Also, in this work, the operators do not need to have dense domains and/or ranges. This generality is not always assumed everywhere since some researchers studied almost sectorial operators defined over domains that are dense as well, see e.g. [2]. Almost sectorial operators frequently appear by the consideration of elliptic operators in regular spaces. Let us now recall briefly some classical examples in this setting. For instance, the minus Laplacian in a bounded domain $\Omega$ is sectorial (this means $\gamma=1$ in Definition 2.1) under some suitable boundary conditions in $L^{p}(\Omega)$ [4, Section 1.3]. Moreover, it is also sectorial in the spaces of bounded or continuous functions [6, 17]. While, in the space of Hölder continuous, functions the minus Laplacian is almost sectorial, see e.g. [19] or [9, Example 3.1.33]. Some other good examples of almost sectorial operators can be found in [14, Section 2]. There is an interesting example of an operator who involves Riemann-Liouville fractional derivatives in the Banach space $C_{\infty}(\mathbb{R};\mathbb{C}^{2})$ of continuous functions which vanish at infinity [15, Example 3.2], that is not sectorial but it is almost sectorial. In [15], more examples of almost sectorial operators are given by using Riemann-Liouville derivatives.
@@
In this paper, we show explicitly the unique solution operator for the abstract Cauchy problem involving a time-dependent variable coefficient and a fractional power of an almost sectorial operator. By using the representation of the solution operator, we then consider two direct abstract Cauchy problems with the same time-variable coefficient, with the target of finding the time-dependent coefficient (inverse abstract Cauchy problems). Here, we mainly follow some ideas from [12] and [14].
@@
We complement our studies, by considering the case of non-autonomous evolution equations with almost sectorial operators. Here, the operator depends on a time-variable. For some works in this direction, we recommend to check e.g. [3, 8]. Again, for this case, we find the unique solution operator (explicitly) for the abstract Cauchy problem with a time-dependent variable coefficient and an almost sectorial operator $A(\tau)$ for $\tau\geqslant 0$. Also, the inverse abstract Cauchy problem with this type of operators is studied.
@@
The structure of the paper is as follows. In Section 2, we recall some definitions and results about almost sectorial operators and abstract evolution equations. We continue with Section 3, where we give some of the main results on the representation of the solution operator of abstract Cauchy problems with time-variable coefficients. While, in Section 4, we consider the inverse abstract Cauchy problem in finding the time-dependent variable coefficient. In the end (Section 5), we study similar problems as mentioned before by means of almost sectorial operators that depends on a time-variable.
@@ 2. Preliminaries
Everywhere below, we assume that $(X,\|\cdot\|)$ is a complex Banach space. Here, an operator in $X$ means a linear map $A:D(A)\subset X\to X$ whose domain $D(A)$ is a linear subspace of $X.$ Also, $\sigma(A)$ is the spectrum of $A,$ and $R(A)$ is the range. Moreover, $\mathscr{L}(X)$ denotes the space of all bounded linear operators on $X.$
Let $0<\mu<\pi$. So, we define by $S_{\mu}^{0}$ to be the open sector $\{z\in\mathbb{C}\setminus\{0\}:|\text{arg}\,z|<\mu\},$ and its closure $S_{\mu}:=\{z\in\mathbb{C}\setminus\{0\}:|\text{arg}\,z|\leqslant\mu\}\cup\{0\}.$ We consider the function arg with values in $(-\pi,\pi].$
@@
Let $-1<\gamma<0$ and $0\leqslant\omega<\pi.$ By $\Theta_{\omega}^{\gamma}(X)$ we denote the set of all closed linear operators $A:D(A)\subset X\to X$ which satisfy
@@
For simplicity we denote $\Theta_{\omega}^{\gamma}$ instead of $\Theta_{\omega}^{\gamma}(X)$. A linear operator $A$ will be called an almost sectorial operator in $X$ if $A\in\Theta_{\omega}^{\gamma}.$ Note that $0\in\rho(A)$ for any $A\in\Theta_{\omega}^{\gamma}.$ Also, we have that $A$ is injective [14, Remark 2.2]. It is important to recall that operators in the class $\Theta_{\omega}^{\gamma}$ have the possibility of having non-dense domain and/or range. This latter feature gives a different panorama with respect to classical results where it is always assumed dense domains [13, 16].
Let us now recall some necessary definitions and results that will be used through the whole paper.
Below we always denote by $\Gamma_{\theta}$ $(0<\theta<\pi)$ the path
@@
First, we take $0<\alpha<\frac{\pi}{2\omega}$ and $t\in S_{\frac{\pi}{2}-\alpha\omega}^{0}$, with $\omega<\mu<\text{min}\left\{\pi,\frac{\pi-2|\text{arg}\,t|}{2\alpha}\right\}.$ Suppose that the function $g_{\alpha,t}(z)=e^{-tz^{\alpha}}$ is defined on $z\in S_{\mu}^{0}.$ Assume also that $A\in\Theta_{\omega}^{\gamma}.$ So, by [14, Lemma 2.13], we have that
$\mathscr{T}_{\alpha}(t)=\frac{1}{2\pi}\int_{\Gamma_{\theta}}e^{-tz^{\alpha}}(z-A)^{-1}{\rm d}z,\quad\omega<\theta<\mu,$
is a bounded linear operator in $X.$ In particular, the above holds for $0<\alpha\leqslant 1$ Note that the above expression for $\mathscr{T}_{\alpha}(t)$ could be also denoted by $e^{-tA^{\alpha}}.$ The latter notation is frequently used for strongly continuous semigroups generated by $A^{\alpha}$.
Next the definition about the analytic semigroups of growth order.
@@
Let $20<\mu<\pi/$ 2 and $\kappa>0.$ A family $\{\mathscr{T}(t):t\in S_{\mu}^{0}\}$ is said to be an analytic semigroup of growth order $\kappa$ if the following conditions hold:
  1. (1)
@@
The above definition has two difference with respect to the one given by Da Prato. First, the set $X_{0}=\bigcup_{t>0}\mathscr{T}(t)X$ not need to be dense in $X.$ Second, the strong continuity of the mapping $t\to\mathscr{T}(t)$ for $t>0$ is replaced by condition (2).
Now we remember some properties of the operators families $\{\mathscr{T}_{\alpha}(t):t\in S_{\frac{\pi}{2}-\alpha\omega}\}$ stated in [14, Theorem 3.9] that will be used implicitly in some of our proofs.
@@
Suppose that $A\in\Theta_{\omega}^{\gamma}$ for some $0<\omega<\frac{\pi}{2}$ and $0<\alpha<\frac{\pi}{2\omega}.$ Then the family $\{\mathscr{T}_{\alpha}(t):t\in S_{\frac{\pi}{2}-\alpha\omega}^{0}\}$ is an analytic semigroup of growth order $\frac{1+\gamma}{\alpha}.$ So, the following assertions are true.
@@
    The function $t\to\mathscr{T}_{\alpha}(t)$ is analytic in $S_{\frac{\pi}{2}-\alpha\omega}^{0}$ and
@@
Abstract evolution equations. Suppose that $A\in\Theta_{\omega}^{\gamma}$ with $20<\omega<\pi/$ 2 and $-1<\gamma<0.$ As usual with this type of operators, we do not request density of the domain of $A.$
Consider the following abstract Cauchy problem:
(2.1) $\begin{split}\partial_{t}u:=u_{t}(t)&=A^{\alpha}u(t),\quad 0<t<T,\quad 0<\alpha<\frac{\pi}{2\omega},\\ u(0)&=u_{0}\in X.\end{split}$
By a classical solution of problem (2.1), we mean a function $u\in C\big{(}(0,T);X\big{)}\cap C^{1}\big{(}(0,T);X\big{)}$ which takes values in $D(A^{\alpha})$ for any $0<t<T$ and satisfies (2.1). It is important to highlight that if $u_{0}\in D(A^{\alpha})$ then the classical solution is continuous also in $t=0$ whenever $\alpha>1+\gamma$ (Theorem 2.3, item (5)).
It can be proved (see e.g. [14, Chapter 4]) that equation (2.1) has a unique classical solution given by
@@ 3. Direct abstract Cauchy problem
In this section, we discuss the direct abstract Cauchy problem with a time-dependent variable coefficient $\phi$ by using an almost sectorial operator in the class $\Theta_{\omega}^{\gamma}$. The time-dependent function $\phi:[0,+\infty)\to[0,+\infty)$ is continuous with $\phi(t)>0$ for $t>0$.
We consider the abstract Cauchy problem in a complex Banach space $X$ with the coefficient $\phi(t)$ and $A\in\Theta_{\omega}^{\gamma}$ $\big{(}0<\omega<\pi/2\big{)}$ as follows:
(3.1) $\begin{split}u_{t}(t)&=\phi(t)A^{\alpha}u(t),\quad 0<t<T,\quad 0<\alpha<\frac{\pi}{2\omega},\\ u(0)&=u_{0}\in X.\end{split}$
@@
Since $\phi$ is a continuous function such that $\phi(t)>0$ for any $t>0$, we have that $\eta(t)=\int_{0}^{t}\phi(r){\rm d}r$ for $t\geqslant 0$ is a bijection. Hence, there exits a unique function $\zeta:[0,+\infty)\to[0,+\infty)$ such that
(3.3) $\begin{split}\zeta(\eta(t))&=t,\quad t\geqslant 0,\\ \eta(\zeta(s))&=s,\quad s\geqslant 0.\end{split}$
Moreover, we get that $\partial_{s}\zeta(s)=\frac{1}{\phi(\zeta(s))}$ for any $s\geqslant 0$. So, we compute
$\begin{split}\frac{\rm d}{{\rm d}s}\big{(}u(\zeta(s))\big{)}&=\partial_{s}u(\zeta(s))\partial_{s}\zeta(s)=\big{(}\phi(\zeta(s))A^{\alpha}u(\zeta(s))\big{)}\frac{1}{\phi(\zeta(s))}=A^{\alpha}u(\zeta(s)),\end{split}$
where $u(\zeta(0))=u(0)=u_{0},$ for any $u$ satisfying equation (3.1). By using the uniqueness of the solution (2.2) of problem (2.1), we obtain
(3.4) $(u\circ\zeta)(s)=\mathscr{T}_{\alpha}(s)u_{0}\quad\text{for any}\quad s\geqslant 0.$
With $s=\eta(t)$, expression in (3.4) can be rewritten as
$u(t)=u(\zeta(\eta(t)))=\mathscr{T}_{\alpha}\big{(}\eta(t)\big{)}u_{0}=\mathscr{T}_{\alpha}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0},$
where $u(t)$ is the solution operator of (3.1). ∎
@@
(3.5) $\begin{split}u_{t}(t)&=\phi(t)Au(t),\quad 0<t<T,\\ u(0)&=u_{0}\in X,\end{split}$
where $A\in\Theta_{\omega}^{\gamma}$ with $20<\omega<\pi/$ 2 . Below we denote by $\mathscr{T}(t),$ $\mathscr{T}_{\phi}(t)$, the analytic semigroup $\mathscr{T}_{1}(t)$, and $\mathscr{T}_{1,\phi}(t)$ respectively.
@@
(3.6) $\mathscr{T}_{\phi}(t)=\mathscr{T}\left\{\left(\int_{0}^{t}\phi(s){\rm d}s\right)\right\},\quad 0<t<T.$

### v2

@@ Abstract.
We derive the explicit solution operator of an abstract Cauchy problem involving a time-variable coefficient and a fractional power of an almost sectorial operator. The time-variable coefficient is recovered by solving the inverse abstract Cauchy problem using the solution operator representation. As a complement, we also study similar problems by considering almost sectorial operators that depend on a time-variable.
@@
The paper has been accepted for publication in the journal “Fractional Calculus and Applied Analysis”.
@@ 1. Introduction
In functional analysis, specifically in operator theory, the Hille–Yosida theorem characterizes the generators of strongly continuous semigroups of linear operators on Banach spaces. That is, for a linear and densely defined operator $A:D(A)\subset X\to X$ ($X$ a complex Banach space), the following estimate holds:
$\|(\lambda I-A)^{-n}\|\leqslant\frac{C}{(\lambda-\omega)^{n}},\quad\lambda>\omega,$
@@
for all positive integers $n$ and any $\lambda$ in the resolvent set $\rho(A).$ For contraction semigroups, the Hille–Yosida theorem can be easily characterized by those operators $A$ such that $\|(\lambda I-A)^{-1}\|\leqslant\frac{1}{\lambda},$ for any $\lambda\in\rho(A)$ with $\lambda>0.$ Nevertheless, there are different classes of important and interesting operators that do not satisfy the estimates above. We recall the so-called almost sectorial operators, see Definition 2.1. More details about such operators can be found e.g. in [12, 16, 20]. It is known that operators satisfying the condition (b) do not generate a $C_{0}$-semigroup since they do not satisfy the Hille–Yosida estimates. But, in this case, the operator $A$ will generate a type of semigroup called an analytic semigroup of growth order $\gamma$ (see Definition 2.3). This concept was first introduced by Da Prato in [5] for positive integer orders. The generalization for any positive order was given in different works e.g. [12, 13, 20, 23] and the references therein. These semigroups are not strongly continuous at $t=0,$ and this is one of the main differences compared with the $C_{0}$-semigroups. Also, in this work, the operators do not need to have dense domains and/or ranges. This generality is not always assumed everywhere since some researchers studied almost sectorial operators defined over domains that are dense as well, see e.g. [2]. Almost sectorial operators frequently appear by the consideration of elliptic operators in regular spaces. Let us briefly recall some classical examples in this setting. For instance, the negative Laplacian in a bounded domain $\Omega$ is sectorial (this means $\gamma=1$ in Definition 2.1) under some suitable boundary conditions in $L^{p}(\Omega)$ [4, Section 1.3]. Moreover, it is also sectorial in spaces of bounded or continuous functions [6, 7, 19]. While, in the space of Hölder continuous functions a translation of the Laplacian is almost sectorial, see e.g. [21] or [11, Example 3.1.33]. Some other good examples of almost sectorial operators can be found in [16, Section 2]. An example of an almost sectorial operator that is not sectorial and involves Riemann-Liouville fractional derivatives in the Banach space $C_{\infty}(\mathbb{R};\mathbb{C}^{2})$ (continuous functions that vanish at infinity) can be found in [17, Example 3.2]. In [17], more examples of almost sectorial operators are given using Riemann-Liouville derivatives.
@@
In this paper, we show explicitly the unique solution operator for the abstract Cauchy problem involving a time-dependent variable coefficient and a fractional power of an almost sectorial operator. By using the representation of the solution operator, we then consider two direct abstract Cauchy problems with the same time-variable coefficient, with the target of finding the time-dependent coefficient (inverse abstract Cauchy problems). To the best of our knowledge, it seems that it is the first time that the latter problems have been considered by means of almost sectorial operators. Here, we mainly follow some ideas from [14] and [16]. Note that the functional calculus developed for almost sectorial operators in [16] will help us to construct the solution operators of our equations. This approach is totally different from the one used in [14]. Also, in this paper, we consider ordinary time derivatives; but future works will be studied with time-fractional derivatives [9, 10].
@@
We complement our studies, by considering the case of non-autonomous evolution equations with almost sectorial operators. Here, the operator depends on a time-variable. For some works in this direction, we recommend checking e.g. [3, 8]. Again, for this case, we find the unique solution operator (explicitly) for the abstract Cauchy problem with a time-dependent variable coefficient and an almost sectorial operator $A(\tau)$ for $\tau\geqslant 0$. We also study an inverse abstract Cauchy problem.
@@
The structure of the paper is as follows. In Section 2, we recall some definitions and results on almost sectorial operators and abstract evolution equations. We continue with Section 3, where we give some of the main results on the representation of the solution operator of abstract Cauchy problems with time-variable coefficients. Section 4 is devoted to the study of the inverse abstract Cauchy problem in finding the time-dependent variable coefficient. In the last Section 5, we also study non-autonomous direct and inverse Cauchy problems. In this case, almost sectorial operators depend on a time-variable.
@@ 2. Preliminaries
We begin by recalling some definitions and results that will be used throughout the paper. We always assume that $(X,\|\cdot\|)$ is a complex Banach space. An operator in $X$ means a linear map $A:D(A)\subset X\to X$ whose domain $D(A)$ is a linear subspace of $X.$ We denote by $\sigma(A)$ the spectrum of $A,$ and by $R(A)$ its range. For $0<\mu<\pi,$ we denote by $S_{\mu}^{0}$, the open sector $\{z\in\mathbb{C}\setminus\{0\}:|\text{arg}\,z|<\mu\},$ and its closure $S_{\mu}:=\{z\in\mathbb{C}\setminus\{0\}:|\text{arg}\,z|\leqslant\mu\}\cup\{0\}.$ We consider the function arg with values in $(-\pi,\pi].$ Set
$\mathcal{F}_{0}^{\gamma}(S_{\mu}^{0})=\bigcup_{s<0}\Psi_{s}^{\gamma}(S_{\mu}^{0})\cup\Psi_{0}(S_{\mu}^{0}),$
and
$\mathcal{F}(S_{\mu}^{0})=\{f\in\mathcal{H}(S_{\mu}^{0}):\,\,\text{there exists}\,\,k,n\in\mathbb{N}\,\,\text{such that}\,\,f\psi_{n}^{k}\in\mathcal{F}_{0}^{\gamma}(S_{\mu}^{0})\},$
where
$\mathcal{H}(S_{\mu}^{0})=\{f:S_{\mu}^{0}\to\mathbb{C};\,\,f\,\,\text{is holomorphic}\},$
$\mathcal{H}^{\infty}(S_{\mu}^{0})=\{f\in\mathcal{H}(S_{\mu}^{0}),\,\,f\,\,\text{is bounded}\},$
$\varphi_{0}(z)=\frac{1}{1+z},\quad\psi_{n}(z)=\frac{z}{(1+z)^{n}},\quad z\in\mathbb{C}\setminus\{-1\},\,\,n\in\mathbb{N}\cup\{0\},$
$\Psi_{0}(S_{\mu}^{0})=\left\{f\in\mathcal{H}(S_{\mu}^{0}):\sup_{z\in S_{\mu}^{0}}\left|\frac{f(z)}{\varphi_{0}(z)}\right|<+\infty\right\},$
and for each $s<0,$
$\Psi_{s}^{\gamma}(S_{\mu}^{0})=\left\{f\in\mathcal{H}(S_{\mu}^{0}):\sup_{z\in S_{\mu}^{0}}|\psi_{n}^{s}(z)f(z)|<+\infty\right\},$
where $n$ is the smallest integer such that $n\geqslant 2$ and $\gamma+1<-(n+1)s.$ It follows
$\mathcal{F}_{0}^{\gamma}(S_{\mu}^{0})\subset\mathcal{H}^{\infty}(S_{\mu}^{0})\subset\mathcal{F}(S_{\mu}^{0})\subset\mathcal{H}(S_{\mu}^{0}),$
and for $k,n\in\mathbb{N}\cup\{0\}$ with $n>k$, one has $\psi_{n}^{k}\in\mathcal{F}_{0}^{\gamma}(S_{\mu}^{0}).$
@@
Let $1<\gamma<0$ and $0\leqslant\omega<\pi.$ By $\Theta_{\omega}^{\gamma}(X)$ we denote the set of all closed linear operators $A:D(A)\subset X\to X$ which satisfy
@@
For simplicity, we write $\Theta_{\omega}^{\gamma}$ instead of $\Theta_{\omega}^{\gamma}(X)$. A linear operator $A$ will be called an almost sectorial operator in $X$ if $A\in\Theta_{\omega}^{\gamma}.$ Note that $0\in\rho(A)$ for any $A\in\Theta_{\omega}^{\gamma}.$ Also, $A$ is injective [16, Remark 2.2]. The operators in $\Theta_{\omega}^{\gamma}$ have the possibility of having non-dense domain and/or range. This latter feature gives a different panorama with respect to the classical results, where dense domains are assumed [15, 18].
We usually denote by $\Gamma_{\theta}$ $(0<\theta<\pi)$ the path
@@
Let $A\in\Theta_{\omega}^{\gamma},$ $\omega<\theta<\mu<\pi,$ and $f\in\mathcal{F}(S_{\mu}^{0}).$ For $k,m\in\mathbb{N}$ such that $f\psi_{m}^{k}\in\mathcal{F}_{0}^{\gamma}(S_{\mu}^{0}),$ we define the linear operator $f(A)$ in the domain $D(f(A))=\{x\in X:(f\psi_{m}^{k})(A)x\in D(A^{(m-1)k})\}$ by
$f(A)=(\psi_{m}^{k}(A))^{-1}\frac{1}{2\pi i}\int_{\Gamma_{\theta}}(f\psi_{m}^{k})(z)(z-A)^{-1}{\rm d}z,$
where the above integral is absolutely convergent and defines a bounded linear operator on $X.$
For every $\beta\in\mathbb{C},$ the function $h_{\beta}(z)=z^{\beta}$ for all $z\in\mathbb{C}\setminus(-\infty,0]$ belongs to the class $\mathcal{F}(S_{\nu}^{0})$ for $0<\nu<\pi,$ and hence we can define $A^{\beta}=h_{\beta}(A).$ Some of the main properties of $A^{\beta}$ ($A\in\Theta_{\omega}^{\gamma},$$\beta\in\mathbb{C}$) are the following [16, Theorem 3.2]:
  1. (i)
    The operator $A^{\beta}$ is closed.
  2. (ii)
    $A^{\beta}A^{\beta^{*}}\subset A^{\beta+\beta^{*}},$ for $\beta^{*}\in\mathbb{C}.$ If $D(A^{\beta+\beta^{*}})\subset D(A^{\beta^{*}}),$ then $A^{\beta}A^{\beta^{*}}=A^{\beta+\beta^{*}}.$
  3. (iii)
    $A^{\beta}$ is injective and $(A^{\beta})^{-1}=A^{-\beta}.$
Take $0<\alpha<\frac{\pi}{2\omega}$ and $t\in S_{\frac{\pi}{2}-\alpha\omega}^{0}$, with $\omega<\mu<\text{min}\left\{\pi,\frac{\pi-2|\text{arg}\,t|}{2\alpha}\right\}.$ Suppose that the function $g_{\alpha,t}(z)=e^{-tz^{\alpha}}$ is defined on $z\in S_{\mu}^{0}.$ Assume also that $A\in\Theta_{\omega}^{\gamma}.$ From [16, Lemma 2.13], we conclude that
(2.1) $\mathscr{T}_{\alpha}(t)=\frac{1}{2\pi}\int_{\Gamma_{\theta}}e^{-tz^{\alpha}}(z-A)^{-1}{\rm d}z,\quad\omega<\theta<\mu,$
is a bounded linear operator in $X.$ In particular, the above holds for $0<\alpha\leqslant 1.$ Note that the above expression for $\mathscr{T}_{\alpha}(t)$ could also be denoted by $e^{-tA^{\alpha}}.$ The latter notation is frequently used for strongly continuous semigroups generated by $A^{\alpha}$. For almost sectorial operators, $A^{\alpha}$ does not always generate a strongly continuous semigroup, see item (5) of Theorem 2.4. This semigroup is singular at $t=0,$ see item (2) of Theorem 2.4.
The following definition was introduced in [5]. Analytic semigroups of growth order are defined. The notion is adapted from the $C_{0}$- semigroups due to the semigroup property holding for these kinds of operators.
Definition 2.3.
Let $0<\mu<\pi/2$ and $\kappa>0.$ A family $\{\mathscr{T}(t):t\in S_{\mu}^{0}\}$ is said to be an analytic semigroup of growth order $\kappa$ if the following conditions hold:
  1. (I)
@@
The above definition has two differences with respect to the one given by Da Prato. First, the set $X_{0}=\bigcup_{t>0}\mathscr{T}(t)X$ is not necessarily dense in $X.$ Second, the strong continuity of the mapping $t\to\mathscr{T}(t)$ for $t>0$ is replaced by condition (II).
We recall some properties of the families $\{\mathscr{T}_{\alpha}(t):t\in S_{\frac{\pi}{2}-\alpha\omega}\}$ stated in [16, Theorem 3.9] that will be used implicitly in some of our proofs.
@@
Suppose that $A\in\Theta_{\omega}^{\gamma}$ for some $0<\omega<\frac{\pi}{2}$ and $0<\alpha<\frac{\pi}{2\omega}.$ Then the family $\{\mathscr{T}_{\alpha}(t):t\in S_{\frac{\pi}{2}-\alpha\omega}^{0}\}$ is an analytic semigroup of growth order $\frac{1+\gamma}{\alpha}.$ So the following assertions are true.
@@
    The function $t\mapsto\mathscr{T}_{\alpha}(t)$ is analytic in $S_{\frac{\pi}{2}-\alpha\omega}^{0}$ and
@@
Remark 2.5.
It is important to highlight and stress that, in general, the analytic semigroups of some growth order are not strongly continuous at $t=0,$ see e.g. Definition 2.3 and Theorem 2.4 (item 5).
Abstract evolution equations. Suppose that $A\in\Theta_{\omega}^{\gamma}$ with $0<\omega<\pi/2$ and $1<\gamma<0.$ Consider the following abstract Cauchy problem:
(2.2) $\begin{split}\partial_{t}u(t):=u_{t}(t)&=A^{\alpha}u(t),\quad 0<t<T,\quad 0<\alpha<\frac{\pi}{2\omega},\\ u(0)&=u_{0}\in D(A^{\alpha}).\end{split}$
By a classical solution of problem (2.2), we mean a function
$u\in C\big([0,T);X\big)\cap C^{1}\big((0,T);X\big)$
which takes values in $D(A^{\alpha})$ for any $0<t<T$ and satisfies (2.2). It is important to note that if $u_{0}\in D(A^{\alpha})$ and $\alpha>1+\gamma$ then there exists a classical solution of (2.2) (Theorem 2.4, item (5)). It can be proved (see e.g. [16, Chapter 4]) that equation (2.2) has a unique classical solution given by
@@ 3. Direct abstract Cauchy problem
In this section, we study the direct abstract Cauchy problem with a time-dependent variable coefficient $\phi$ and an almost sectorial operator. We consider the abstract Cauchy problem in a complex Banach space $X$ with the coefficient $\phi(t)$ and $A\in\Theta_{\omega}^{\gamma}$ $\big(0<\omega<\pi/2\big)$ as follows:
(3.1) $\begin{split}u_{t}(t)&=\phi(t)A^{\alpha}u(t),\quad 0<t<T,\quad 1+\gamma<\alpha<\frac{\pi}{2\omega},\\ u(0)&=u_{0}\in D(A^{\alpha}),\end{split}$
where the time-dependent function $\phi:[0,+\infty)\to[0,+\infty)$ is continuous such that $\phi(t)>0$ for $t>0.$
@@
Since $\phi(t)>0$ for $t>0$ and (2.1), it follows
$\mathscr{T}_{\alpha,\phi}(t)=\frac{1}{2\pi}\int_{\Gamma_{\theta}}e^{-z^{\alpha}\int_{0}^{t}\phi(s){\rm d}s}(z-A)^{-1}{\rm d}z,\quad\omega<\theta<\mu<\text{min}\left\{\pi,\frac{\pi}{2\alpha}\right\},$
is a bounded linear operator in $X.$ Now, by Theorem 2.4 (item (4)), we obtain
$\frac{{\rm d}}{{\rm d}t}\mathscr{T}_{\alpha,\phi}(t)=\phi(t)A^{\alpha}\mathscr{T}_{\alpha,\phi}(t).$
From Theorem 2.4 (item (5)) and $\alpha>1+\gamma$, $u(t)=\mathscr{T}_{\alpha,\phi}(t)u_{0}$ is a classical solution of problem (3.1). ∎
@@
where $A\in\Theta_{\omega}^{\gamma}$ with $0<\omega<\pi/2.$ In what follows, we denote by $\mathscr{T}(t),$ $\mathscr{T}_{\phi}(t)$, the operators $\mathscr{T}_{1}(t)$, and $\mathscr{T}_{1,\phi}(t)$ respectively.
@@
$\mathscr{T}_{\phi}(t)=\mathscr{T}\left\{\left(\int_{0}^{t}\phi(s){\rm d}s\right)\right\},\quad 0<t<T.$